OpenStudy (anonymous):

The sum of the first 20 terms of the sequence a sub n = 7n + 2 is _____. Confused... any explanation will be appreciated :)

jimthompson5910 (jim_thompson5910):

What's the first term?

OpenStudy (anonymous):

2 right?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

what's the 20th term?

jimthompson5910 (jim_thompson5910):

oh wait, are you starting at n = 0 or n = 1?

OpenStudy (anonymous):

is that 20*2?

OpenStudy (anonymous):

see thats the confusing part idk if its 1 or 0

jimthompson5910 (jim_thompson5910):

hmm maybe your book will tell you which one to start with because both are valid

OpenStudy (anonymous):

i think its 0

jimthompson5910 (jim_thompson5910):

Alright, let's go with that then

jimthompson5910 (jim_thompson5910):

So if n = 0, then a0 = 2 is the first term n = 19 is the 20th term (since we started at 0 and not 1)

jimthompson5910 (jim_thompson5910):

I mean, you plug in n = 19 to find the 20th term

OpenStudy (anonymous):

so 7*19?

OpenStudy (anonymous):

and if it was 1 it would be 7*20?

jimthompson5910 (jim_thompson5910):

7*19+2

OpenStudy (anonymous):

oh ok and then thats the answer

jimthompson5910 (jim_thompson5910):

no, you need to find the first and 20th terms to help get you to the answer

jimthompson5910 (jim_thompson5910):

The formula is Sum of first n terms = (n*(first term + nth term))/2

jimthompson5910 (jim_thompson5910):

So Sum of first 20 terms = (20*(first term + 20th term))/2

OpenStudy (anonymous):

is the first term 2

jimthompson5910 (jim_thompson5910):

yes, if you start at n = 0

OpenStudy (anonymous):

so Sum of first 20 terms = (20*(2 + 20th term))/2

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

so then (40+20th term)/2??

jimthompson5910 (jim_thompson5910):

first off, what is the 20th term? that would be very helpful

jimthompson5910 (jim_thompson5910):

remember, plug in n = 19 and evaluate

OpenStudy (anonymous):

oh its135

OpenStudy (anonymous):

(40+135)/2

jimthompson5910 (jim_thompson5910):

so it's really Sum of first 20 terms = (20*(2 + 135))/2

jimthompson5910 (jim_thompson5910):

Sum of first 20 terms = (20*(2 + 135))/2 Sum of first 20 terms = (20*(137))/2 Sum of first 20 terms = (2740)/2 Sum of first 20 terms = 1370 This is of course assuming we start at n = 0

OpenStudy (anonymous):

oh ok i see 1370

OpenStudy (anonymous):

Thank you soo much @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yw